# STBA: Towards Evaluating the Robustness of DNNs for Query-Limited Black-box Scenario

### v1

@@ III-D Optimization
Recall that our goal is to calculate the flow field $\bm{f}$, which will be applied to the clean image’s high-frequency part $\bm{x}_{high}$ to generate its adversarial counterpart one $\bm{x}_{high}^{adv}$ and further formulate adversarial example $\bm{x}^{adv}$. we regard the flow field $\bm{f}$ comes from a unique distribution and re-defined the expected adversarial example $\bm{x}^{adv}$ by re-written the Eq.III-B as
@@
TABLE I: The victim models and their corresponding classification accuracy (%) on the validation set of CIFAR-10 and CIFAR-100, and the subset of ImageNet dataset NIPS2017, respectively.
@@
TABLE II: Attack success rate (ASR,%) against normal models on three benchmark datasets under query budget is set to 1000 and 10000.
@@
Dataset Model Square RS_Attack ASH NPAttack STBA Square RS_Attack ASH NPAttack STBA
CIFAR-10 VGG-19 90.27 88.99 96.47 90.91 99.64 99.65 91.38 96.51 99.70 100.00
ResNet-56 97.14 62.79 98.44 94.31 100.00 99.94 67.74 98.44 100.00 100.00
MobileNetv2 98.99 82.32 98.74 97.10 99.97 99.77 62.49 98.75 100.00 100.00
ShuffleNetv2 99.56 84.18 96.83 92.01 99.86 99.89 58.13 96.83 100.00 100.00
CIFAR-100 VGG-19 92.93 90.48 95.12 88.31 98.17 98.83 93.12 95.12 98.66 99.98
ResNet-56 99.82 84.18 98.93 95.47 100.00 99.86 86.88 98.93 99.97 100.00
MobileNetv2 99.47 82.32 98.51 96.34 100.00 99.80 84.08 98.54 99.74 100.00
ShuffleNetv2 97.83 68.79 94.78 90.61 99.77 99.84 71.25 94.78 99.79 100.00
STL-10 VGG-19 89.25 42.97 95.12 95.07 99.00 96.57 56.32 91.05 99.43 100.00
ResNet-56 96.78 28.12 91.41 90.34 99.80 100.00 42.99 91.57 100.00 100.00
MobileNetv2 98.36 33.59 91.99 96.66 97.10 99.85 64.24 89.65 100.00 100.00
ShuffleNetv2 96.82 33.13 90.82 94.75 98.50 100.00 42.15 90.43 100.00 100.00
ImageNet VGG-19 81.25 42.97 95.12 95.07 99.52 99.48 92.36 96.09 100.00 100.00
ResNet-152 96.78 28.12 91.41 90.48 95.87 98.61 78.82 93.36 99.84 100.00
MobileNetv2 98.36 33.59 91.99 96.66 99.36 100.00 83.16 92.38 100.00 100.00
ShuffleNetv2 96.82 33.13 90.82 94.75 99.05 99.29 82.47 91.99 100.00 100.00
TABLE III: Attack success rate (ASR,%) against robust models on three benchmark datasets under query budget is set to 1000 and 10000.
@@
Dataset Model Square RS_Attack ASH NPAttack STBA Square RS_Attack ASH NPAttack STBA
CIFAR-10 HYDRA 13.33 29.87 28.99 7.49 67.30 20.86 31.56 29.18 19.68 98.24
Wangadv 13.32 24.10 22.13 8.79 61.53 20.67 25.12 22.42 20.38 93.27
Zhangadv 18.45 37.37 36.57 10.99 78.67 26.35 38.61 36.77 27.37 98.14
FastAdv 23.14 37.27 43.26 13.79 80.12 34.05 38.67 43.33 32.97 96.83
CIFAR-100 Pangadv 26.38 51.23 40.25 13.39 82.17 33.89 52.82 40.25 27.07 98.34
Efficientadv 31.60 40.14 39.83 12.99 83.83 41.96 41.96 39.83 30.48 98.52
Proxy 27.78 51.84 39.34 11.09 84.84 53.47 53.47 39.62 27.17 98.87
Learnadv 28.16 53.88 48.21 9.99 95.93 55.03 55.03 49.12 28.97 99.58
STL-10 Free_AT 54.75 46.35 79.04 25.87 61.58 64.06 46.35 83.66 50.19 97.34
FastAdv 47.80 35.85 73.31 25.77 63.20 60.95 35.85 79.95 51.36 97.92
TRADES 60.52 24.78 80.21 51.25 90.93 82.01 24.78 85.09 92.22 100.00
MART 53.23 35.85 73.24 48.55 88.97 72.59 29.61 80.08 82.88 100.00
ImageNet Slman_${R18}$ 39.26 36.33 49.02 36.33 70.02 52.73 75.35 67.38 59.53 97.53
Slman_${R50}$ 17.97 18.55 36.13 18.55 54.84 33.79 64.58 55.08 43.97 95.45
Engsrom_${R18}$ 18.36 18.55 40.82 18.55 44.97 38.67 65.62 62.11 46.69 95.64
FastAdv 36.72 30.86 55.86 30.86 50.47 55.47 76.39 72.27 53.70 96.20
@@ IV-A Settings
To investigate the performance of the proposed method in attacking robust models, we select some of the most recent defense techniques as follows: they are including HYDRA [53], Wangadv [54], Zhangadv [55], FastAdv [56] for CIFAR-10, Pangadv [57], Efficientadv [58], Proxy [59], Learnadv [60] for CIFAR-100, FreeAT [61], FastAdv [56], TRADES [55] and MART [54] for STL-10, andSalmanR50 [62], Engstromadv [63], FastAdv [56], SalmanR18 [62] for ImageNet. For all the models of CIFAR-10, CIFAR-100, and ImageNet, we use their implementation in the robustbench toolbox [^7] https://github.com/RobustBench/robustbench [64] and the models’ weights are also provided in this toolbox [64].For all these models, we chose their $L_{\infty}$-norm version parameters because we carry out $L_{\infty}$-norm attacks in this paper. The same as normal models, we train the robust models of STL-10 by ourselves with the help of their officially provided codes, and we listed the recognition performance of all these aforementioned models on tested images in the Table. I.
@@
Baselines: The baselines for query-limited score-based black-box attack settings are Square Attack [1], RS_Attack [2], ASH [3] and NPAttack [4]. For Square Attack, ASH and NPAttack, we set the $L_{\infty}=0.05$, and for RS_Attack, we use its patches version called RS-Attack and set the patch size as 3x3 for CIFAR-10 and CIFAR-100 datasets, 8x8 for STL-10, and 20x20 for ImageNet dataset.
Metrics: We compare our proposed method with several query-efficient black-box attack techniques concerned with Attack Success Rate (ASR) and image quality involving LPIPS [65], DISTS [66], MAD [67], BRISQUE [68], FID [69], SSIM [70], PSNR, VIF [71] and FSIM [72].
Figure 5: The ASR matrix of baselines and the proposed STBA on CIFAR-10, STL-10 and ImageNet, respectively. The row represents which model is targeted during the generation of adversarial examples, while the column represents which model is attacked by the generated examples.
@@ IV-B Quantitative Comparison with the Query-Efficient Black-box Attacks
In this subsection, we will evaluate the proposed STBA and the baselines in attacking performance and query efficiency on the validation dataset of CIFAR-10 and CIFAR-100, and the whole ImageNet dataset.
@@
The empirical results indicated that although previous techniques show fantastic performance in query-based black-box attack settings, once we put a relatively extreme limit on the maximum number of queries, these methods will lose their advantages and show dissatisfaction results. Besides, such results also verify our assumption that the previous techniques struggle to attack the robustly trained model successfully.
@@
Figure 7: The clean example and their corresponding adversarial noise generated by baselines and the proposed STBA. The first row is the clean images, and the flowing rows are the adversarial noise generated by Square Attack [1], ASH [3], NPAttack [4] and STBA, respectively.
@@ IV-C Transferability
The motivation of the current work states that the generation of adversarial examples is generally based on the assumption of transferability, which means that an adversarial example generated according to a model can be used to attack other unseen models. To see if this assumption is valid in black-box settings, here, we follow the previous works [19, 73] to examine the transferability of the generated adversarial examples across different models on CIFAR-10, STL-10 and ImageNet. Specifically, we generate adversarial examples for the aforementioned eight models for each dataset, including four normal-trained and four adversarial-trained. Following the settings in [74], we randomly select 1000 images from the CIFAR-10 and STL-10 test set, which are classified correctly by the model, whereas the corresponding adversarial examples are misclassified; for ImageNet (NIPS2017), we selected all the correctly classified and the corresponding adversarial examples are misclassified ones for this experiment. Where the generated adversarial examples are used to attack the other models. Here, we set the maximal query number to 1000 for all baselines.
We illustrate the results of this part with the ASR confusion matrix on the three datasets, as shown in Fig. 5. The row represents which model is targeted during the generation of adversarial examples, while the column represents which model is attacked by the generated examples. From this figure, we can see that the highest transferability ASR on three datasets of STBA can achieve 74.7%, 57.8%, and 73.2%, respectively. In contrast, the comparable baseline method NPAttack is 66.5%, 47.7% and 26.6%, respectively. The rest of the baselines even exhibit poor transferability. It means that the examples generated by STBA produce higher transfer ASR in most cases than baselines, especially for the STL-10 and ImageNet datasets, which validate the superior transferability of STBA. Meanwhile, this phenomenon shows that baselines heavily rely on the feedback of the target model during each query and cannot extract transferable features. By contrast, our method learns the adversarial flow field $\bm{f}$ that does not collapse to a specific model.
In addition, we are surprised to find that adversarial examples generated on the robust model have better transfer attack performance, which means that the researchers can generate adversarial examples on the robust model and then implement transfer attacks on other models to explore existing models’ vulnerability more effectively and further promote the defenders to propose more effective defense mechanisms to enhance the models’ robustness.
@@ IV-D Visual Quality
The quantitative experimental results of images’ quality can be seen in Table. IV, which indicated that the proposed method has the lowest LPIPS, DISTS, MAD, BRISQUE, and FID (the lower is better), are 0.0162, 0.0680, 3.0883, 16.482, and 13.9358, respectively, and has the highest SSIM, PSNR, VIF and FSIM (the higher is better), achieving 0.9879, 37.3324, 0.8258 and 0.9873, respectively, in comparison to the baselines on ImageNet dataset. The results show that the proposed method is superior to the existing attack methods.
TABLE IV: Visual quality evaluation results calculated on adversarial examples crafted by Square, ASH, NPAttack, and the proposed STBA.
Metric Square ASH NPAttack STBA
LPIPS $\downarrow$ 0.2448 0.1298 0.1168 0.0162
DISTS $\downarrow$ 0.2677 0.1625 0.1340 0.0680
MAD $\downarrow$ 90.7684 73.7850 66.6841 3.0883
BRISQUE $\downarrow$ 29.7395 20.3779 17.2939 16.482
FID $\downarrow$ 97.0623 50.4029 29.9010 13.9358
SSIM $\uparrow$ 0.8202 0.9516 0.9122 0.9879
PSNR $\uparrow$ 26.1883 26.1909 31.3421 37.3324
VIF $\uparrow$ 0.6240 0.7887 0.5404 0.8258
FSIM $\uparrow$ 0.9025 0.9630 0.9410 0.9873
Furthermore, to better observe the difference between the adversarial examples generated by our method and the baselines from the visual aspect, we draw the adversarial perturbation generated on ImageNet by baselines and the proposed method in Fig. 7, the target model is pre-trained VGG-16. The first row is the benign examples, and the following are the adversarial noise of Square [1], ASH [3], NPAttack [4] and our method, respectively. Noted that, for better observation, we magnified the noise by multiplying by a factor of 25. From Fig. 7, we can clearly observe that baseline methods distort the image without ordering. In contrast, the adversarial perturbations generated by our method are focused on the target object instead of disturbing the whole image, and its noise contains more semantic information and is more imperceptible to human eyes.

### v2

@@ III-D Optimization
Recall that our goal is to calculate the flow field $\bm{f}$, which will be applied to the clean image’s high-frequency part $\bm{x}_{high}$ to generate its adversarial counterpart one $\bm{x}_{high}^{adv}$ and further formulate adversarial example $\bm{x}^{adv}$. we regard the flow field $\bm{f}$ comes from a unique distribution and re-defined the expected adversarial example $\bm{x}^{adv}$ by re-written the Eq.4 and Eq.5 as
@@
Algorithm 1 Spatial Transform for Query-Efficient Attack
Parameter: The flow field $\bm{f}=[0,1]^{2*h*w}$.
Output: The adversarial example $\bm{x}^{adv}$ used for attack.
1:  Initialize the flow budget $\xi$ with $\xi=10^{-1}$.
2:  Initialize $\bm{\mu}$ randomly.
3:  Initialize the parameters of the flow $\bm{f_{0}}$ with zeros.
4:  Decompose $\bm{x}$ to $\bm{x}_{high}$ and $\bm{x}_{low}$.
5:  while $q<=\bm{Q}_{max}/n_{sample}$ do
7:     Set $\bm{f}_{k}=\bm{\delta_{f_{k}}}$ for all $k=1,...,n_{sample}$.
8:     Compute ${\bm{x}_{high}^{adv}}^{k}$ by Eq. 14 for all $k=1,...,n_{sample}$.
[…]
11:     Compute $\nabla_{\bm{\mu}}\bm{J}(\bm{\mu},\sigma)=\frac{1}{n_{sample}}{\textstyle\sum_{k=1}^{n_{sample}}}\hat{\mathcal{L}}\cdot\bm{\epsilon}_{k}$.
12:     update $\bm{\mu}\longleftarrow\bm{\mu}-lr\nabla_{\bm{\mu}}\bm{J}(\bm{\mu},\sigma)$.
13:     if $q\ \%\ adjust_{num}=0$ then
14:        Update $\xi_{t}$ by Eq.8.
15:     end if
17:     if $\bm{x}^{adv}_{t}$ attack $\bm{\mathcal{C}}$ successfully then
18:        break.
19:     end if
20:  end while
21:  return $\bm{x}^{adv}_{t}$
TABLE I: The victim models and their corresponding classification accuracy (%) on the validation set of CIFAR-10, CIFAR-100, STL-10 and the subset of ImageNet dataset NIPS2017, respectively.
@@
TABLE II: Attack success rate (ASR,%) against normal models on four benchmark datasets under query budget is set to 1000 and 10000
@@
Dataset Model Square AdvFlow RS ASH NP Dif STBA Square AdvFlow RS ASH NP Dif STBA
CIFAR-10 VGG-19 90.27 64.71 88.99 96.47 90.91 96.89 99.64 99.65 97.05 91.38 96.51 99.70 99.97 100.00
ResNet-56 97.14 69.19 62.79 98.44 94.31 99.13 100.00 99.94 99.64 67.74 98.44 100.00 100.00 100.00
MobileNetv2 98.99 79.46 82.32 98.74 97.10 99.58 99.97 99.77 99.74 62.49 98.75 100.00 100.00 100.00
ShuffleNetv2 99.56 65.78 84.18 96.83 92.01 97.12 99.86 99.89 98.80 58.13 96.83 100.00 100.00 100.00
CIFAR-100 VGG-19 92.93 64.37 90.48 95.12 88.31 91.58 98.17 98.83 96.03 93.12 95.12 98.66 99.42 99.98
ResNet-56 99.82 90.32 84.18 98.93 95.47 99.06 100.00 99.86 99.93 86.88 98.93 99.97 100.00 100.00
MobileNetv2 99.47 88.83 82.32 98.51 96.34 98.92 100.00 99.80 99.82 84.08 98.54 99.74 100.00 100.00
ShuffleNetv2 97.83 77.76 68.79 94.78 90.61 95.80 99.77 99.84 98.99 71.25 94.78 99.79 99.96 100.00
STL-10 VGG-19 89.25 77.70 42.97 95.12 95.07 83.65 99.00 96.57 99.47 56.32 91.05 99.43 99.75 100.00
ResNet-56 96.78 73.94 28.12 91.41 90.34 96.90 99.80 100.00 99.69 42.99 91.57 100.00 100.00 100.00
MobileNetv2 98.36 77.32 33.59 91.99 96.66 91.95 97.10 99.85 99.40 64.24 89.65 100.00 100.00 100.00
ShuffleNetv2 96.82 79.08 33.13 90.82 94.75 92.17 98.50 100.00 99.81 42.15 90.43 100.00 99.79 100.00
ImageNet VGG-19 81.25 67.09 42.97 95.12 95.07 93.48 99.52 99.48 99.21 92.36 96.09 100.00 99.84 100.00
ResNet-152 96.78 64.07 28.12 91.41 90.48 82.19 95.87 98.61 99.52 78.82 93.36 99.84 99.68 100.00
MobileNetv2 98.36 79.01 33.59 91.99 96.66 95.55 99.36 100.00 99.84 83.16 92.38 100.00 100.00 100.00
ShuffleNetv2 96.82 79.01 33.13 90.82 94.75 89.19 99.05 99.29 99.84 82.47 91.99 100.00 99.84 100.00
TABLE III: Attack success rate (ASR,%) against robust models on four benchmark datasets under query budget is set to 1000 and 10000.
@@
Dataset Model Square AdvFlow RS ASH NP Dif STBA Square AdvFlow RS ASH NP Dif STBA
CIFAR-10 HYDRA 13.33 36.24 29.87 28.99 7.49 21.89 67.30 20.86 58.60 31.56 29.18 19.68 47.19 98.24
Wangadv 13.32 33.58 24.10 22.13 8.79 20.79 61.53 20.67 54.82 25.12 22.42 20.38 43.50 93.27
Zhangadv 18.45 43.17 37.37 36.57 10.99 27.86 78.67 26.35 63.84 38.61 36.77 27.37 52.22 98.14
FastAdv 23.14 50.39 37.27 43.26 13.79 33.49 80.12 34.05 72.88 38.67 43.33 32.97 61.57 96.83
CIFAR-100 Pangadv 26.38 38.35 51.23 40.25 13.39 30.47 82.17 33.89 61.38 52.82 40.25 27.07 56.41 98.34
Efficientadv 31.60 43.11 40.14 39.83 12.99 35.27 83.83 41.96 68.81 41.96 39.83 30.48 100.00 98.52
Proxy 27.78 39.14 51.84 39.34 11.09 31.36 84.84 53.47 64.46 53.47 39.62 27.17 99.06 98.87
Learnadv 28.16 36.42 53.88 48.21 9.99 32.69 95.93 55.03 67.78 55.03 49.12 28.97 87.76 99.58
STL-10 Free_AT 54.75 75.07 46.35 79.04 25.87 33.35 61.58 64.06 91.24 46.35 83.66 50.19 70.40 97.34
FastAdv 47.80 73.99 35.85 73.31 25.77 33.86 63.20 60.95 92.49 35.85 79.95 51.36 71.43 97.92
TRADES 60.52 69.49 24.78 80.21 51.25 54.75 90.93 82.01 94.93 24.78 85.09 92.22 93.57 100.00
MART 53.23 71.73 35.85 73.24 48.55 49.69 88.97 72.59 92.60 29.61 80.08 82.88 87.02 100.00
ImageNet Slman_${R18}$ 39.26 31.12 36.33 49.02 36.33 8.16 70.02 52.73 70.78 75.35 67.38 59.53 70.78 97.53
Slman_${R50}$ 17.97 34.91 18.55 36.13 18.55 8.35 54.84 33.79 70.97 64.58 55.08 43.97 70.97 95.45
Engsrom_${R18}$ 18.36 36.81 18.55 40.82 18.55 9.49 44.97 38.67 73.06 65.62 62.11 46.69 73.06 95.64
FastAdv 36.72 51.80 30.86 55.86 30.86 18.41 50.47 55.47 86.15 76.39 72.27 53.70 86.15 96.20
@@ IV-A Settings
To investigate the performance of the proposed method in attacking robust models, we select some of the most recent defense techniques as follows: they are including HYDRA [55], Wangadv [56], Zhangadv [57], FastAdv [58] for CIFAR-10, Pangadv [59], Efficientadv [60], Proxy [61], Learnadv [62] for CIFAR-100, FreeAT [63], FastAdv [58], TRADES [57] and MART [56] for STL-10, and SalmanR50 [64], Engstromadv [65], FastAdv [58], SalmanR18 [64] for ImageNet. For all the models of CIFAR-10, CIFAR-100, and ImageNet, we use their implementation in the robustbench toolbox [^7] https://github.com/RobustBench/robustbench [66] and the models’ weights are also provided in this toolbox [66]. For all these models, we chose their $L_{\infty}$-norm version parameters because we carry out $L_{\infty}$-norm attacks in this paper. Similar to normal models, we train the robust models of STL-10 by ourselves with the help of their officially provided codes, and we listed the recognition performance of all these aforementioned models on tested images in the Table. I.
@@
Baselines: The baselines for query-limited score-based black-box attack settings are Square Attack [1], AdvFlow [2], RS_Attack [3], ASH [4], NPAttack [5] and DifAttack [6]. For Square Attack, AdvFlow, ASH, and NPAttack, we set the $L_{\infty}=0.05$, and for RS_Attack, we set the patch size as 3x3 for CIFAR-10 and CIFAR-100, 8x8 for STL-10, and 20x20 for ImageNet. For AdvFlow, we use the officially pre-trained flow weights on CIFAR-10 and train on other datasets by ourselves, and for the DifAttack, we use the SqueezeNet [67] as the surrogate model.
Metrics: We compare our proposed method with several query-efficient black-box attack techniques concerned with Attack Success Rate (ASR), Average Query number (Average.Q) and Median Query number (Med.Q), and image quality involving LPIPS [68], DISTS [69], MAD [70], BRISQUE [71], FID [72], SSIM [73], PSNR, VIF [74] and FSIM [75].
Figure 5: Confusion matrix of transferability for adversarial attacks generated by baselines and the proposed STBA on CIFAR-10, CIFAR-100, STL-10 and ImageNet, respectively. The row represents which model is targeted during the generation of adversarial examples, while the column represents which model is attacked by such synthesized examples.
@@ IV-B Quantitative Comparison with the Query-Efficient Black-box Attacks
In this subsection, we will evaluate the proposed STBA and the baselines in attacking performance and query efficiency on the validation dataset of CIFAR-10, CIFAR-100, STL-10 and the whole NIPS2017 dataset. Note that, in this part, we use RS, NP, and Dif to represent RS_Attack, NPAttack, and DifAttack, respectively.
@@
The empirical results indicated that although previous techniques show fantastic performance in query-based black-box attack settings, once we put a relatively extreme limit on the maximum number of queries, these methods will lose their advantages and show dissatisfactory results. Besides, such results also verify our assumption that the previous techniques struggle to attack the robustly trained model successfully.
@@
Figure 7: The clean example and their corresponding adversarial noise generated by baselines and the proposed STBA. The first row is the clean images, and the flowing rows are the adversarial noise generated by Square Attack [1], AdvFlow [2], ASH [4], NPAttack [5], DifAttack [6], and STBA, respectively.
@@ IV-C Transferability
The motivation of the current work states that the generation of adversarial examples is generally based on the assumption of transferability, which means that an adversarial example generated according to a model can be used to attack other unseen models. To see if this assumption is valid in black-box settings, here, we follow the previous works [2, 76] to examine the transferability of the generated adversarial examples across different models on CIFAR-10, CIFAR-100, STL-10 and ImageNet. Specifically, we generate adversarial examples for the aforementioned eight models shown in Table I for each dataset, including four normal-trained and four adversarial-trained. We randomly selected 1000 images from the CIFAR-10, CIFAR-10, and STL-10 test set, and selected all images from ImageNet (NIPS2017), which are classified correctly by these models, whereas the corresponding adversarial examples are misclassified, for this experiment. Where such generated adversarial examples are used to attack the other models. Here, we set the maximal query number to 1000 for all baselines.
We illustrate the results of this part with the ASR confusion matrix on these four datasets, as shown in Fig. 5. The row represents which model is targeted during the generation of adversarial examples, while the column represents which model is attacked by the generated examples. From this figure, we can see that the highest transferability ASR on four datasets of STBA can achieve 74%, 83%, 57%, and 73%, respectively. In contrast, the comparable baseline method NPAttack is 66%, 84%, 47%, and 26%, respectively. The rest of the baselines even exhibit poor transferability. It means that the examples generated by STBA produce higher transfer ASR in most cases than baselines, especially for the STL-10 and ImageNet datasets, which validate the superior transferability of STBA. Meanwhile, this phenomenon shows that baselines heavily rely on the feedback of the target model during each query and cannot extract transferable features. By contrast, our method learns the adversarial flow field $\bm{f}$ that does not collapse to a specific model.
In addition, we were intrigued to discover that AEs generated on robust models exhibit superior transfer attack performance. This finding suggests that researchers can effectively leverage AEs crafted on robust models to conduct transfer attacks on various other models. Such a strategy allows for a more efficient exploration of vulnerabilities across existing models. Furthermore, this approach can significantly aid in the identification of critical weaknesses, thereby encouraging defenders to develop more robust and effective defense mechanisms. This could lead to enhancements in overall model robustness, promoting more secure AI applications.
@@ IV-D Visual Quality
The quantitative experimental results of images’ quality can be seen in Table. IV, which indicated that the proposed method has the lowest MAD and BRISQUE (the lower is better), are 3.0883 and 16.4823, respectively, and has the highest SSIM, PSNR, VIF and FSIM (the higher is better), achieving 0.9879, 37.3324, 0.8258 and 0.9873, respectively, in comparison to the baselines on ImageNet dataset. The results show that the proposed method is superior to the existing attack methods.
TABLE IV: Visual quality evaluation results calculated on adversarial examples crafted by Square, AdvFlow, ASH, NPAttack, DifAttack, and the proposed STBA.
Metric Square AdvFlow ASH NPAttack DifAttack STBA
LPIPS $\downarrow$ 0.2448 0.0135 0.1298 0.1168 0.2079 0.0162
DISTS $\downarrow$ 0.2677 0.0366 0.1625 0.1340 0.1869 0.0680
MAD $\downarrow$ 90.7684 29.4512 73.7850 66.6841 87.9537 3.0883
BRISQUE $\downarrow$ 29.7395 17.0994 20.3779 17.2939 26.6368 16.4823
FID $\downarrow$ 97.0623 8.5572 50.4029 29.9010 44.4464 13.9358
SSIM $\uparrow$ 0.8202 0.9865 0.9516 0.9122 0.7969 0.9879
PSNR $\uparrow$ 26.1883 33.3259 26.1909 31.3421 28.1149 37.3324
VIF $\uparrow$ 0.6240 0.8137 0.7887 0.5404 0.4456 0.8258
FSIM $\uparrow$ 0.9025 0.9721 0.9630 0.9410 0.9001 0.9873
Furthermore, to better observe the difference between the adversarial examples generated by our method and the baselines from the visual aspect, we draw the adversarial perturbation generated on ImageNet by baselines and the proposed method in Fig. 7, the target model is pre-trained VGG-16. The first row is the benign examples, and the following are the adversarial noise of Square [1], AdvFlow [2], ASH [4], NPAttack [5], DifAttack [6] and our method, respectively. Noted that, for better observation, we magnified the noise by multiplying a factor of 25. From Fig. 7, we can clearly observe that baseline methods distort the image without ordering. In contrast, the adversarial perturbations generated by our method are focused on the target object instead of disturbing the whole image, and its noise contains more semantic information and is more imperceptible to human eyes.